in a right angled triangle if one of the acute angle is twicethe other,prove that the hypotenuse is double the side opposite to smallest angle.

Answer by cleomenius(959)

x + 2x + 90 = 180
3x = 90
x = 30 degrees, the smallest angle.
sin(30) = opp/hyp
1/2 = x/hyp
hyp = 2 * x.
Cleomenius.